# USTC-ICTS/PCFT-26-06

### v1

@@ V.2.2 Quadratic action for scalar perturbation
The next step is to construct a single scalar field dark energy model that is capable of crossing the cosmological constant boundary. It is well known that quintessence or phantom models cannot achieve such a crossing. Hence a more general scalar field Lagrangian must be considered. In this paper, we adopt a single scalar field model with a generalized kinetic term together with an additional noncanonical kinetic term. The Lagrangian under consideration is therefore given by:
$\mathcal{L}(X,\phi_{1})=c_{1}X+c_{2}\sqrt{X}-V(\phi_{1}),$ (57)
where $c_{1}$ and $c_{2}$ are assumed to be non-zero constants. The potential $V(\phi_{1})$ is chosen to have the following form:
@@
$3H^{2}=\rho_{m}+\rho_{{}_{DE}},\quad-2\dot{H}-H^{2}=P_{m}+P_{{}_{DE}},\quad c_{1}\ddot{\phi}_{1}+3H\left(c_{1}\dot{\phi}_{1}+\frac{c_{2}}{\sqrt{2}}\mathcal{P}\right)+V_{\phi_{1}}=0,$ (59)
@@
$\rho_{{}_{DE}}=\frac{c_{1}}{2}\dot{\phi}_{1}^{\,2}+V\left(\phi_{1}\right),\quad P_{{}_{DE}}=\frac{c_{1}}{2}\dot{\phi}_{1}^{\,2}+\frac{c_{2}}{\sqrt{2}}|\dot{\phi}_{1}|-V\left(\phi_{1}\right).$ (60)
@@
Since the energy density $\rho_{{}_{DE}}$ is positive definite during the evolution of $\phi_{1}$, the denominator of the EoS parameter in Eq.(61) must remain positive during the evolution. Meanwhile, if $c_{1}$ and $c_{2}$ have the same sign, the numerator cannot cross zero. Therefore, a sign difference, $c_{1}c_{2}<0$, is generally required for the crossing. For simplicity, in the numerical examples presented below we set $c_{1}>0$, $c_{2}<0$.
@@
We next numerically solve the background equations in (59). To illustrate the evolution of the EoS parameter $w_{{}_{DE}}$, we have assumed $\Omega_{m,0}\approx 0.31$ and $h\approx 0.69$. In addition, the rescaling parameter $M\approx 8.77\times 10^{-61}$ is also introduced to simplify the calculations. Within the above parameter ranges, the model parameters can be chosen with considerable flexibility. In particular, the evolution of the EoS $w_{{}_{DE}}$ as a function of redshift $z$ is presented in FIG.4(b), where we adopt the parameter values: $c_{1}=0.8,~c_{2}/M=-1.2,~\Lambda_{\phi}/M^{2}=0.66,~m_{\phi}/M=1.14$ and $\lambda_{\phi}/M^{2}=0.8\,$. The initial conditions are fixed to $\dot{\phi}_{1_{ini}}/M=-1.06$ and $\phi_{1_{ini}}=0.7$. In this case, the EoS parameter $w_{{}_{DE}}(z)$ is found to evolve smoothly and crosses the cosmological constant boundary at $z\approx 0.43$. At the present epoch, i.e., $z=0$, the model yields $w_{0}\approx-0.8$. In addition, the phase space trajectory of the solution for this model is also plotted in FIG.4(b). One can see directly that $|\dot{\phi}_{1}|<1$ during the evolution. Therefore, under these conditions the quartic term in $V(\phi_{1})$ is subdominant to the quadratic term.
@@
(b) Left panel: evolution of EoS parameter $w_{{}_{DE}}(z)$ as a function of redshift $z$ and the gray dashed line marks the cosmological constant $\Lambda$ with EoS parameter w=1w=-1. Right panel: the phase space trajectory of the solution with initial conditions $\dot{\phi}_{1_{ini}}/M=-1.06$ and $\phi_{1_{ini}}=0.7$.
In addition to the results presented above, the data set is used to explore the effect of varying a single parameter, while keeping the remaining parameters fixed, on the evolution of $w_{{}_{DE}}(z)$. After some calculations, the results are displayed in FIG.4(c). In the upper left panel $w_{{}_{DE}}(z)$ is shown for a range of values of $c_{1}$: as $c_{1}$ is increased, the redshift of cosmological constant boundary crossing shifts to lower $z$, and the present day value $w_{0}$ moves closer to -1. The dependence on $c_{2}$ is shown in the upper right panel: smaller values of $c_{2}$ are associated with a larger crossing redshift and a less negative $w_{0}$. The lower left panel illustrates the effect of $\Lambda_{\phi}$: upon decreasing $\Lambda_{\phi}$ the crossing redshift is reduced and $w_{0}$ approaches closer to -1. The impact of the parameter $m_{\phi}$ is presented in the lower right panel. The larger $m_{\phi}$ will lead to a smaller crossing redshift and a larger value of present day $w_{0}$. Since the effect on $w_{{}_{DE}}$ of increasing $\lambda_{\phi}$ is similar to that of decreasing $m_{\phi}$, results for variations of $\lambda_{\phi}$ are not shown here.
(c) The way $w_{{}_{DE}}(z)$ changes with different values of a certain parameter. Upper left panel corresponds to different values of $c_{1}$ and upper right panel is associated with different values of $c_{2}/M$. Lower left panel is related to different values of $\Lambda_{\phi}/M^{2}$ and lower right panel is associated with different values of $m_{\phi}/M$. The gray dashed line shows the EoS parameter w=1w=-1, which is corresponding to the cosmological constant $\Lambda$.
VI Conclusion
@@
Current observational data favor a dynamical dark energy with quintom-B behavior. However, in Einstein gravity with minimal couplings, the unavoidable instabilities in perturbations prevent either a single perfect fluid or a single kessence-like scalar field from realizing the quintom scenario. In this paper we proposed a possible way out of this difficulty by incorporating the dark energy into the NYTG model, which is a slight modification to the GR equivalent teleparallel gravity. In our model, the quintom dark energy is still represented by a single perfect fluid or a single kessence-like scalar field, but it has a direct coupling to the Nieh-Yan density. Through the analysis on the cosmological perturbation theory of this model, we showed that the introduced coupling brings an extra constraint and forces the gauge-invariant dark energy perturbation to vanish. In this way, the dark energy perturbation is removed from the menu of the dynamical degrees of freedom and the original instabilities in it are also eliminated. Overall, our work provides a new approach to the investigation of quintom dark energy models.

### v2

@@ V.2.2 Quadratic action for scalar perturbation
The next step is to construct a single scalar field dark energy model that is capable of crossing the cosmological constant boundary. It is well known that quintessence or phantom models cannot achieve such a crossing. Hence a more general scalar field Lagrangian must be considered. In this paper, we extended the quintessence model with an additional noncanonical kinetic term. The Lagrangian under consideration is therefore given by:
$\mathcal{L}(X,\phi_{1})=X+c_{1}\sqrt{X}-V(\phi_{1}),$ (57)
where $c_{1}$ is assumed to be non-zero constant. The potential $V(\phi_{1})$ is chosen to have the following form:
@@
$3H^{2}=\rho_{m}+\rho_{{}_{DE}},\quad-2\dot{H}-H^{2}=P_{m}+P_{{}_{DE}},\quad\ddot{\phi}_{1}+3H\left(\dot{\phi}_{1}+\frac{c_{1}}{\sqrt{2}}\mathcal{P}\right)+V_{\phi_{1}}=0,$ (59)
@@
$\rho_{{}_{DE}}=\frac{\dot{\phi}_{1}^{\,2}}{2}+V\left(\phi_{1}\right),\quad P_{{}_{DE}}=\frac{\dot{\phi}_{1}^{\,2}}{2}+\frac{c_{1}}{\sqrt{2}}|\dot{\phi}_{1}|-V\left(\phi_{1}\right).$ (60)
@@
Since the energy density $\rho_{{}_{DE}}$ is positive definite during the evolution of $\phi_{1}$, the denominator of the EoS parameter in Eq.(61) must remain positive during the evolution. Meanwhile, if $c_{1}$ is positive, the numerator cannot cross zero. Therefore, a negative $c_{1}$ is generally required for the crossing.
@@
We next numerically solve the background equations in (59). To illustrate the evolution of the EoS parameter $w_{{}_{DE}}$, we have assumed $\Omega_{m,0}\approx 0.31$ and $h\approx 0.69$. Within the above parameter ranges, the model parameters can be chosen with considerable flexibility. In particular, the evolution of the EoS $w_{{}_{DE}}$ as a function of redshift $z$ is presented in FIG.3(a), where we adopt the parameter values: $c_{1}=-1.14\times 10^{-60},~\Lambda_{\phi}=5.29\times 10^{-121},~m_{\phi}=1.12\times 10^{-60}$ and $\lambda_{\phi}=7.69\times 10^{-121}\,$. The initial conditions are fixed to $\dot{\phi}_{1_{ini}}=-8.06\times 10^{-61}$ and $\phi_{1_{ini}}=0.6$. In this case, the EoS parameter $w_{{}_{DE}}(z)$ is found to evolve smoothly and crosses the cosmological constant boundary at $z\approx 0.44$. At the present epoch, i.e., $z=0$, the model yields $w_{0}\approx-0.8$. We emphasize that the model considered here serves only as an illustrative example, and a detailed comparison with observational data will be pursued in future work.
@@
(a) The evolution of EoS parameter $w_{{}_{DE}}(z)$ as a function of redshift $z$ and the gray dashed line marks the cosmological constant $\Lambda$ with EoS parameter w=1w=-1.
VI Parity violation in gravitational waves
In this section we demonstrate that our model not only avoid the instabilities in dark energy perturbation, but also gives rise to parity violation in gravitational waves. From the linear perturbation equations one can find that the tensor $N^{\mu\nu}$ in Eq.(20) does not contain any vector perturbation Li et al. (2020). Hence the vector sector contains no dynamical degrees of freedom, as in GR. We next consider the tensor sector, where tensor perturbations arise only from the tetrad $e^{i}_{\phantom{1}j}$:
$e^{i}_{\phantom{1}j}=a\left(\delta_{ij}+\frac{1}{2}h^{T}_{ij}\right).$ (62)
The tensor perturbations $h^{T}_{ij}$ are transverse and traceless, i.e., $\partial_{i}h^{T}_{ij}=\delta^{ij}h^{T}_{ij}=0$. No matter dark energy is a single perfect fluid or single kessence-like scalar field, the quadratic action for tensor perturbations always takes the same form, which coincides with that obtained in Ref.Li et al. (2020):
$S=\int d^{4}x\ \frac{a^{2}}{8}\left({h^{T}_{ij}}^{\prime}{h^{T}_{ij}}^{\prime}-\partial_{k}h^{T}_{ij}\partial_{k}h^{T}_{ij}-2c\phi_{1}^{\prime}\epsilon_{ijk}h^{T}_{il}\partial_{j}h^{T}_{kl}\right)\,$, (63)
After expanding the tensor perturbations in Fourier space by circular polarization bases
$h_{ij}^{T}(t,\vec{x})=\frac{1}{(2\pi)^{3/2}}\int d^{3}\vec{k}~h_{ij}^{T}(t,\vec{k})\,e^{i\vec{k}\cdot\vec{x}}~,~h_{ij}^{T}(t,\vec{k})=\sum_{A=L,R}h_{A}(t,\vec{k})\,\hat{e}^{A}_{ij}(\vec{k}),$ (64)
the quadratic action can be rewritten as
$S=\sum_{A=L,R}\int d\eta d^{3}k\ \frac{a^{2}}{4}\left[h^{\prime\,2}_{A}-(k^{2}+2\mathfrak{p}_{A}c\phi_{1}^{\prime}k)h^{2}_{A}\right]\,$, (65)
where $\mathfrak{p}_{L}=-1,\ \mathfrak{p}_{R}=1$, and we mark $h^{\prime*}_{A}h^{\prime}_{A}\rightarrow h^{\prime 2}_{A},\ h^{*}_{A}h_{A}\rightarrow h^{2}_{A}$ for simplicity. The equation for tensor perturbations derived from (65) is
$h_{A}^{\prime\prime}+2\mathcal{H}h_{A}^{\prime}+\omega_{A}^{2}h_{A}=0,$ (66)
where $\omega_{A}^{2}=k^{2}+2\mathfrak{p}_{A}c\phi_{1}^{\prime}k$. From (66), one can find that gravitational waves with different helicities will have different phase velocities: $v_{p}^{A}=\omega_{A}/k\approx 1+\mathfrak{p}_{A}c\phi_{1}^{\prime}/k\equiv 1+\mathfrak{p}_{A}aM_{PV}/(2k)$. This phenomenon is known as velocity birefringence of gravitational waves and provides an explicit signal of parity violation in this model. This is a unique theoretical prediction of the mechanism proposed in this paper, makes it distinguishable from other models of non-minimal couplings or modified gravities. It has obtained constraints by current experimental data of gravitational waves and will accept more tests in the future. Through the full Bayesian inference on the gravitational wave events of binary black hole merges (BBH), an upper bound on the parameters of NYTG model was placed in Ref.Wu et al. (2022): $M_{PV}<6.5\times 10^{-42}$GeV. In our framework this implies the constraint: $|c\phi_{1}^{\prime}/a|<3.25\times 10^{-42}$GeV.
VII Conclusion
@@
Current observational data favor a dynamical dark energy with quintom-B behavior. However, in Einstein gravity with minimal couplings, the unavoidable instabilities in perturbations prevent either a single perfect fluid or a single kessence-like scalar field from realizing the quintom scenario. In this paper we proposed a possible way out of this difficulty by incorporating the dark energy into the NYTG model, which is a slight modification to the GR equivalent teleparallel gravity. In our model, the quintom dark energy is still represented by a single perfect fluid or a single kessence-like scalar field, but it has a direct coupling to the Nieh-Yan density. Through the analyses on the cosmological perturbations of this model, we showed that the introduced coupling brings an extra constraint and forces the gauge-invariant dark energy perturbation to vanish. In this way, the dark energy perturbation is removed from the menu of the dynamical degrees of freedom and the original instabilities in it are also eliminated. We also showed that the Nieh-Yan coupling of dark energy leads inevitably parity violations in gravitational waves. This can be considered as unique theoretical prediction of this mechanism, further signals the dynamical nature of dark energy, and is expected to be tested by future gravitational wave experiments.
In summary, our work provides a new approach to the investigation of quintom dark energy models.